# Synthesizing Efficient and Permissive Programmatic Runtime Shields for Neural Policies

### v1

@@ 3.1. Program Sketch
We construct the program sketch as follows:
1def programmatic_runtime_shield($c$, $x_{1}$, $x_{2}$, …, $x_{n}$):
[…]
4 return $\mathbf{K}$
5 else:
6 return $c$
This sketch partially implements a runtime shield that takes the command $c$ from the neural policy and the environment state $x_{1},x_{2},...,x_{n}$ as input and returns a safe command. There is a linear policy that produces a command $\mathbf{K}$ based on the environment state variables (line 2, marked in green), where $\theta_{1},\dots,\theta_{n}$ are coefficients in the linear policy to be synthesized. Another main component is a conditional statement, which is a switching strategy that determines when to invoke the linear policy to replace the unsafe command $c$ with the safe command $\mathbf{K}$ (line 3, marked in red), where $\lambda$ is the threshold in the switching strategy to be synthesized. The sketch is then instantiated by filling the holes with the parameters synthesized by Aegis.
@@ 3.2. Inferring Environment Dynamics
As defined in Section 2, an environment state $s_{t}$ transitions to the subsequent state $s_{t+1}$ based on the state transition function $s_{t+1}=s_{t}+f(s_{t},c_{t})\times\Delta t$, where $c_{t}$ represents the commands issued by the neural policy at time step $t$. What we aim to infer is $f(s_{t},c_{t})$, which describes the dynamics of the physical environment, i.e., how the next state evolves in response to both the commands from the neural policy and the current state. We adopt the assumption that the environment dynamics can be approximated by a discrete-time linear time-invariant system model (Rugh, 1996; Zadeh and Desoer, 2008), which is widely used for modeling the dynamics of physical environments (Menghi et al., 2020; Zhu et al., 2019; Lillicrap et al., 2016). Such a model can be defined as follows:
@@
where $\mathbf{A}$ and $\mathbf{B}$ are two constant matrices. The linear time-invariant system model $f^{\prime}(s_{t},c_{t})$ is a linear approximation of the unknown dynamics function $f(s_{t},c_{t})$ around an equilibrium point (Cooke and Huang, 1996; Ramos, 2005), defined by specific environment states and commands $s^{*}$ and $c^{*}$ such that applying $c^{*}$ to $s^{*}$ results in no change in the environment state over time, i.e., $f(s^{*},c^{*})=0$. In linearization, the equilibrium point is often chosen as the initial environment state $s_{0}$ and the command $c_{0}$ at the beginning of the system operation, i.e., $s^{*}=s_{0}$ and $c^{*}=c_{0}$, so that the system’s behavior can be approximated by a linear model right from the start, simplifying the subsequent synthesis and verification processes. In such a case, the matrices $\mathbf{A}$ and $\mathbf{B}$ in $f^{\prime}(s_{t},c_{t})$ can be derived as the Jacobian matrices (Arrowsmith and Place, 1992), which represent the first-order partial derivatives of $f(s_{t},c_{t})$ with respect to $s_{t}$ and $c_{t}$, respectively. Considering $f(s_{t},c_{t})$ as a vector-valued function with $s_{t}$ and $c_{t}$ as input vectors having $m$ and $n$ dimensions and $f(s_{t},c_{t})$ as a $z$-dimensional output vector, respectively, then the Jacobian matrices $\mathbf{A}$ and $\mathbf{B}$ are $z\times m$ and $z\times n$ matrices, respectively, with the elements defined as:
@@
where $f_{i}(s_{t},c_{t})$ denotes the $i$-th element of the output vector $f(s_{t},c_{t})$, $s_{j}$ and $c_{k}$ are the $j$-th element of the state vector $s_{t}$ and the $k$-th element of the command vector $c_{t}$, respectively. The $\mathbf{A}_{ij}$ and $\mathbf{B}_{ik}$ are the elements of the Jacobian matrices $\mathbf{A}$ and $\mathbf{B}$, respectively.
@@
1
[…]
@@ 3.3. Counterexample-Guided Inductive Synthesis
With the inferred environment dynamics function, we are now prepared to synthesize the linear policy, which is the first major component of the programmatic runtime shield we create. Our goal is to synthesize a linear policy that ensures all outputs adhere to the safety specifications. To achieve this, we introduce a counterexample-guided inductive synthesis algorithm, which is a popular approach for synthesizing programs from specifications (Češka et al., 2019; Solar-Lezama, 2009, 2008). This algorithm iteratively refines a candidate program using counterexamples generated during verification, ensuring that the synthesized program satisfies the desired properties. We tailor this algorithm to synthesize the linear policy for the programmatic runtime shield. Specifically, our algorithm initializes a linear policy using the Linear-Quadratic Regulator (LQR) algorithm (Anderson and Moore, 2007) and iteratively refines it based on counterexamples identified by the maximal output admissible set (Gilbert and Tan, 1991). We introduce our algorithm in the following paragraphs.
The LQR algorithm is a classic and effective method for designing optimal control policies for linear time-invariant systems (Anderson and Moore, 2007). It is widely used in various fields, including robotics (Mason et al., 2016), aerospace (Yang, 2014), and economics (Yaghobipour and Yarahmadi, 2020), due to its simplicity and effectiveness in managing multi-variable control systems. LQR derives control policies by minimizing a quadratic cost function, which is the sum of state and command costs. Specifically, given the state transition function $s_{t+1}=s_{t}+f(s_{t},c_{t})\times\Delta t$, where $f(s_{t},c_{t})=\mathbf{A}s_{t}+\mathbf{B}c_{t}$, the objective of the LQR is to determine the control input that minimizes the following quadratic cost function:
(7) $J=\sum_{t=0}^{\infty}\left(s_{t}^{T}\mathbf{Q}s_{t}+c_{t}^{T}\mathbf{R}c_{t}\right)$
where the term $s_{t}^{T}\mathbf{Q}s_{t}$ represents the state cost, and $c_{t}^{T}\mathbf{R}c_{t}$ represents the command cost, with $s_{t}$ and $c_{t}$ being the state and command at time step $t$, respectively. The intuition behind these costs is that the state cost penalizes deviations of the system’s current state from the desired state, encouraging the system to behave as expected, while the command cost penalizes the control commands, discouraging excessive actions and promoting smooth, efficient control that does not overly strain the system. The LQR algorithm computes the optimal control commands as a linear function of the current states, i.e.,
@@
where $\mathbf{K}$ is the matrix of the parameters in the linear function, and is computed by solving the continuous-time Algebraic Riccati Equation (Anderson and Moore, 2007; Benner et al., 2020):
(9) $\mathbf{A}^{T}\mathbf{P}+\mathbf{PA}-\mathbf{PBR}^{-1}\mathbf{B}^{T}\mathbf{P}+\mathbf{Q}=0$
Once $\mathbf{A}^{P}$ is obtained, the matrix $\mathbf{K}$ can be computed as:
(10) $\mathbf{K}=\mathbf{R}^{-1}\mathbf{B}^{T}\mathbf{P}$
Note that in our case we assume that the state cost matrix $\mathbf{Q}$ and the command cost matrix $\mathbf{R}$ are well-defined and provided by the user, and how to choose them is beyond the scope of this paper. We refer interested readers to (Anderson and Moore, 2007) for more details. Moreover, in our work the above calculation is done with an off-the-shelf solver, i.e., pyMPC (Marcucci, 2024).
@@
[…]
@@
Require: $\mathbf{K}$, $\mathcal{Z}$, $f^{\prime}$, $r$, $lr$, $T$
@@ 3       $s_{0}\leftarrow ce$
4       for time step $\text{time step }t\leftarrow 1$ to $T$ do
@@ 10             $\mathbf{K}\leftarrow\mathbf{K}+lr\times\Delta$
Require: $f^{\prime}$, $\mathbf{K}$, $\mathcal{S}_{0}$, $\lnot\mathcal{S}_{u}$, $r$, $lr$, $T$
@@ 3.4. Switching Strategy
We propose using a Bayesian optimization algorithm to solve this problem. Bayesian optimization is a powerful tool for optimizing black-box functions that are expensive to evaluate (Frazier, 2018; Snoek et al., 2012). It maintains a surrogate model, which approximates the real objective function but at a much lower computational cost. During the optimization process, candidate solutions are evaluated using the surrogate model, and the real objective function is evaluated only when a promising solution is identified. This approach significantly reduces the cost of evaluating candidate solutions, making it particularly well-suited for computationally expensive optimization problems.
@@
where the $log$ function can make the optimization more sensitive to the small number of violations, ensuring that the algorithm focuses on minimizing the number of violations. The adding of 1 in the logarithm and the denominator is to avoid undefined values when the number of violations or interventions is zero. Intuitively, the ideal threshold $\lambda$ derived by the optimization algorithm with this objective function should minimize the number of violations to zero while maximizing the ratio of necessary interventions to total interventions, ensuring that the system remains safe with minimal interventions from the linear policy.
@@
In addition, the acquisition function of the Gaussian process is used to select the next candidate solutions over the current best solution. We follow (Chen et al., 2021) to use expected improvement (Benassi et al., 2011) as the acquisition function, which measures the expected improvement of the surrogate model over the current best solution and selects the candidate solution that maximizes the expected improvement. Formally, the expected improvement is defined as:

### v3

@@ 3.2. Program Sketch
Figure 11. The universal program sketch of a runtime shield synthesized by Aegis.
@@
We construct the program sketch as Figure 11. This sketch partially implements a runtime shield that takes the command $c$ from the neural policy and the environment state $x_{1},x_{2},...,x_{n}$ as input and returns a safe command. There is a linear policy that produces a command $\mathbf{K}$ based on the environment state variables (line 2, marked in red), where $\theta_{1},\dots,\theta_{n}$ are coefficients in the linear policy to be synthesized. Another main component is a conditional statement, which is a switching strategy that determines when to invoke the linear policy to replace the unsafe command $c$ with the safe command $\mathbf{K}$ (line 3, marked in green), where $\lambda$ is the threshold in the switching strategy to be synthesized. The sketch is then instantiated by filling the holes with the parameters synthesized by Aegis.
Here, we formulate the shield using a linear policy not only to align with prior work (Zhu et al., 2019), but also due to its ability to approximate key behaviors of neural policies, efficiency, and compatibility with formal safety analysis. Prior studies have demonstrated that linear policies can capture the core functionalities of more complex controllers (Mania et al., 2018; Zhu et al., 2019; Shi et al., 2025b). For instance, Rajeswaran et al. (Rajeswaran et al., 2017) show that while robots controlled by linear policies may not walk like those using neural policies, they can still stand effectively, achieving critical safety-relevant behaviors. Furthermore, linear policies execute significantly faster (Shi et al., 2025b), which is crucial in runtime shielding, where minimizing overhead is essential to maintaining system responsiveness—especially in safety-critical applications where even minor delays in executing commands can have serious consequences. In addition, linear policies enable the use of easily tractable verification techniques, such as maximal output admissible sets (Gilbert and Tan, 1991) or barrier certificates (Prajna and Jadbabaie, 2004), making them well-suited for integration into our synthesis framework. While exploring more complex controllers is an interesting direction, doing so would necessitate a complete redesign of the synthesis pipeline, which we leave to future work.
@@ 3.3. Inferring Environment Dynamics
As defined in Section 2, an environment state $s_{t}$ transitions to the subsequent state $s_{t+1}$ based on the state transition function $s_{t+1}=s_{t}+f(s_{t},c_{t})\times\Delta t$, where $c_{t}$ represents the commands issued by the neural policy at time step $t$. What we aim to infer is $f(s_{t},c_{t})$, which describes the dynamics of the physical environment, i.e., how the next state evolves in response to both the commands from the neural policy and the current state. We adopt the assumption that the environment dynamics can be approximated by a discrete-time linear time-invariant system model (Rugh, 1993; Zadeh and Desoer, 2008), which is widely used for modeling the dynamics of physical environments (Menghi et al., 2020; Zhu et al., 2019; Lillicrap et al., 2016). Such a model can be defined as follows:
@@
where $\mathbf{A}$ and $\mathbf{B}$ are two constant matrices. The linear time-invariant system model $f^{\prime}(s_{t},c_{t})$ is a linear approximation of the unknown dynamics function $f(s_{t},c_{t})$ around an equilibrium point (Cooke and Huang, 2002; Ramos, 2005), defined by specific environment states and commands $s^{*}$ and $c^{*}$ such that applying $c^{*}$ to $s^{*}$ results in no change in the environment state over time, i.e., $f(s^{*},c^{*})=0$. In linearization, the equilibrium point is often chosen as the initial environment state $s_{0}$ and the command $c_{0}$ at the beginning of the system operation, i.e., $s^{*}=s_{0}$ and $c^{*}=c_{0}$, so that the system’s behavior can be approximated by a linear model right from the start, simplifying the subsequent synthesis and verification processes. In such a case, the matrices $\mathbf{A}$ and $\mathbf{B}$ in $f^{\prime}(s_{t},c_{t})$ can be derived as the Jacobian matrices (Arrowsmith and Place, 1992), which represent the first-order partial derivatives of $f(s_{t},c_{t})$ with respect to $s_{t}$ and $c_{t}$, respectively. Considering $f(s_{t},c_{t})$ as a vector-valued function with $s_{t}$ and $c_{t}$ as input vectors having $m$ and $n$ dimensions and $f(s_{t},c_{t})$ as a $z$-dimensional output vector, respectively, then the Jacobian matrices $\mathbf{A}$ and $\mathbf{B}$ are $z\times m$ and $z\times n$ matrices, respectively, with the elements defined as:
@@
[…]
@@
1 $\mathbf{A}\leftarrow\mathbf{0}$, $\mathbf{B}\leftarrow\mathbf{0}$, $\epsilon\leftarrow\text{random\_perturbation}()$
@@ 3.4. Counterexample-Guided Inductive Synthesis
With the inferred environment dynamics function, we are now prepared to synthesize the linear policy, which is the first major component of the programmatic runtime shield we create. Our goal is to synthesize a linear policy that ensures all outputs adhere to the safety specifications. To achieve this, we introduce a counterexample-guided inductive synthesis algorithm, which is a popular approach for synthesizing programs from specifications (Češka et al., 2019; Solar-Lezama, 2009, 2008). This algorithm iteratively refines a candidate program using counterexamples generated during verification, ensuring that the synthesized program satisfies the desired properties. We tailor this algorithm to synthesize the linear policy for the programmatic runtime shield. Specifically, our algorithm initializes a linear policy using the Linear-Quadratic Regulator (LQR) algorithm (Anderson and Moore, 1990) and iteratively refines it based on counterexamples identified by the maximal output admissible set (Gilbert and Tan, 1991). We introduce our algorithm in the following paragraphs.
The LQR algorithm is a classic and effective method for designing optimal control policies for linear time-invariant systems (Anderson and Moore, 1990). It is widely used in various fields, including robotics (Mason et al., 2016), aerospace (Yang, 2014), and economics (Yaghobipour and Yarahmadi, 2020), due to its simplicity and effectiveness in managing multi-variable control systems. LQR derives control policies by minimizing a quadratic cost function, which is the sum of state and command costs. Specifically, given the state transition function $s_{t+1}=s_{t}+f(s_{t},c_{t})\times\Delta t$, where $f(s_{t},c_{t})=\mathbf{A}s_{t}+\mathbf{B}c_{t}$, the objective of the LQR is to determine the control input that minimizes the following quadratic cost function:
(11) $J=\sum_{t=0}^{\infty}\left(s_{t}^{\top}\mathbf{Q}s_{t}+c_{t}^{\top}\mathbf{R}c_{t}\right)$
where the term $s_{t}^{\top}\mathbf{Q}s_{t}$ represents the state cost, and $c_{t}^{\top}\mathbf{R}c_{t}$ represents the command cost, with $s_{t}$ and $c_{t}$ being the state and command at time step $t$, respectively. The intuition behind these costs is that the state cost penalizes deviations of the system’s current state from the desired state, encouraging the system to behave as expected, while the command cost penalizes the control commands, discouraging excessive actions and promoting smooth, efficient control that does not overly strain the system. The LQR algorithm computes the optimal control commands as a linear function of the current states, i.e.,
@@
where $\mathbf{K}$ is the matrix of the parameters in the linear function, and is computed step by step to derive the discrete-time Algebraic Riccati Equation (Anderson and Moore, 1990; Benner et al., 2020). Firstly, given the above quadratic cost function, we need to define a candidate value function $V_{t}(s_{t})$, which is a quadratic function of the state $s_{t}$:
(13) $V_{t}(s_{t})=s_{t}^{\top}\mathbf{P}s_{t}$
where $\mathbf{P}$ is a positive semi-definite matrix that needs to be determined. We then apply the Bellman principle (Bellman, 1957) to define the value function recursively:
(14) $V_{t}(s_{t})=\min_{c_{t}}\left[s_{t}^{\top}\mathbf{Q}s_{t}+c_{t}^{\top}\mathbf{R}c_{t}+V_{t+1}(s_{t+1})\right]$
Then we substitute the dynamics and the value function:
(15) $V_{t}(s_{t})=\min_{c_{t}}\big[$ $s_{t}^{\top}\mathbf{Q}s_{t}+c_{t}^{\top}\mathbf{R}c_{t}+(s_{t}+\Delta t(\mathbf{A}s_{t}+\mathbf{B}c_{t}))^{\top}\mathbf{P}(s_{t}+\Delta t(\mathbf{A}s_{t}+\mathbf{B}c_{t}))\big]$
Define the closed-form expansion:
(16) $s_{t+1}$ $=s_{t}+\Delta t\mathbf{A}s_{t}+\Delta t\mathbf{B}c_{t}=(\mathbf{I}+\Delta t\mathbf{A})s_{t}+\Delta t\mathbf{B}c_{t}$
Let $\tilde{\mathbf{A}}=\mathbf{I}+\Delta t\mathbf{A}$ and $\tilde{\mathbf{B}}=\Delta t\mathbf{B}$. Then:
(17) $s_{t+1}=\tilde{\mathbf{A}}s_{t}+\tilde{\mathbf{B}}c_{t}$
The value function becomes:
(18) $V_{t}(s_{t})=\min_{c_{t}}\Big[$ $s_{t}^{\top}\mathbf{Q}s_{t}+c_{t}^{\top}\mathbf{R}c_{t}+(\tilde{\mathbf{A}}s_{t}+\tilde{\mathbf{B}}c_{t})^{\top}\mathbf{P}(\tilde{\mathbf{A}}s_{t}+\tilde{\mathbf{B}}c_{t})\Big].$
We then rearrange the above equation as:
(19) $V_{t}(s_{t})=\min_{c_{t}}\Big[$ $s_{t}^{\top}(\mathbf{Q}+\tilde{\mathbf{A}}^{\top}\mathbf{P}\tilde{\mathbf{A}})s_{t}+c_{t}^{\top}(\mathbf{R}+\tilde{\mathbf{B}}^{\top}\mathbf{P}\tilde{\mathbf{B}})c_{t}+2s_{t}^{\top}\tilde{\mathbf{A}}^{\top}\mathbf{P}\tilde{\mathbf{B}}c_{t}\Big]$
Taking the derivative with respect to $c_{t}$ and setting to zero:
(20) $\frac{\partial V_{t}}{\partial c_{t}}=2(\mathbf{R}+\tilde{\mathbf{B}}^{\top}\mathbf{P}\tilde{\mathbf{B}})c_{t}+2\tilde{\mathbf{B}}^{\top}\mathbf{P}\tilde{\mathbf{A}}s_{t}=0$
Then we solve $c_{t}$ which gives the optimal control command:
(21) $c_{t}=-\mathbf{K}s_{t},\quad\text{where}\quad\mathbf{K}=(\mathbf{R}+\tilde{\mathbf{B}}^{\top}\mathbf{P}\tilde{\mathbf{B}})^{-1}\tilde{\mathbf{B}}^{\top}\mathbf{P}\tilde{\mathbf{A}}$
Plugging $c_{t}$ back into the value function, we obtain a new expression for $V_{t}(s_{t})$, from which the updated matrix $\mathbf{P}_{t}$ can be derived recursively:
(22) $\mathbf{P}_{t}=\mathbf{Q}+\tilde{\mathbf{A}}^{\top}\mathbf{P}_{t+1}\tilde{\mathbf{A}}-\tilde{\mathbf{A}}^{\top}\mathbf{P}_{t+1}\tilde{\mathbf{B}}(\mathbf{R}+\tilde{\mathbf{B}}^{\top}\mathbf{P}_{t+1}\tilde{\mathbf{B}})^{-1}\tilde{\mathbf{B}}^{\top}\mathbf{P}_{t+1}\tilde{\mathbf{A}}$
This is the discrete-time algebraic Riccati recursion under Euler-discretized dynamics, which can be solved using value iteration (Bellman, 1957; Sutton, 1992) or other numerical methods like Hamiltonian matrices (Lancaster and Rodman, 1995).
Note that in our case we assume that the state cost matrix $\mathbf{Q}$ and the command cost matrix $\mathbf{R}$ are well-defined and provided by the user, and how to choose them is beyond the scope of this paper. We refer interested readers to (Anderson and Moore, 1990) for more details. Moreover, in our work the above calculation is done with an off-the-shelf solver, i.e., pyMPC (Marcucci, 2024).
@@
$\text{ where }\left.s_{t+1}=s_{t}+f^{\prime}\left(s_{t},-\mathbf{K}s_{t}\right)\times\Delta t\right\}$
@@
Require: $\mathbf{K}$, $\mathcal{Z}$, $f^{\prime}$, $r$, lrlr, $T$
@@ 3  $s_{0}\leftarrow ce$
4  for $\text{time step }t\leftarrow 1$ to $T$ do
@@ 11     $\mathbf{K}\leftarrow\mathbf{K}+lr\times\Delta$
Require: $f^{\prime}$, $\mathbf{K}$, $\mathcal{S}_{0}$, $\lnot\mathcal{S}_{u}$, $r$, lrlr, $T$
@@ 6  $\mathcal{Z}\leftarrow\left\{s_{0}\in\mathcal{S}_{0}\mid s_{0}\notin\mathcal{X}\right\}$
Note that the above algorithm is conditionally complete, meaning that if the maximal output admissible set is computable and the algorithm converges, it will eventually synthesize a linear policy that is verified to be safe, with all counterexamples eliminated. This follows simply from the fact that the termination condition of the iterative process is precisely when the computed MOAS fully covers the set of initial environment states $\mathcal{S}_{0}$, at which point no further counterexamples exist (see line 3).
@@ 3.5. Switching Strategy
We propose using a Bayesian optimization algorithm to solve this problem. Bayesian optimization is a powerful tool for optimizing black-box functions that are expensive to evaluate (Couckuyt et al., 2022; Snoek et al., 2012), which is widely used in various software-related domain such as compiler optimization (Hellsten et al., 2024; Chen et al., 2021a). . It maintains a surrogate model, which approximates the real objective function but at a much lower computational cost. During the optimization process, candidate solutions are evaluated using the surrogate model, and the real objective function is evaluated only when a promising solution is identified. This approach significantly reduces the cost of evaluating candidate solutions, making it particularly well-suited for computationally expensive optimization problems.
@@
where the loglog function can make the optimization more sensitive to the small number of violations, ensuring that the algorithm focuses on minimizing the number of violations. The adding of 1 in the logarithm and the denominator is to avoid undefined values when the number of violations or interventions is zero. Intuitively, the ideal threshold $\lambda$ derived by the optimization algorithm with this objective function should minimize the number of violations to zero while maximizing the ratio of necessary interventions to total interventions, ensuring that the system remains safe with minimal interventions from the linear policy.
@@
In addition, the acquisition function of the Gaussian process is used to select the next candidate solutions over the current best solution. We follow (Chen et al., 2021b) to use expected improvement (Benassi et al., 2011) as the acquisition function, which measures the expected improvement of the surrogate model over the current best solution and selects the candidate solution that maximizes the expected improvement. Formally, the expected improvement is defined as:
@@ 4. Evaluation
4.1. Experimental Setup
Experimental Subjects
To comprehensively evaluate Aegis, we select eight control systems that are widely used in prior work (Zhu et al., 2019; Shi et al., 2025b), covering a range of real-world scenarios including autonomous driving (Self-Driving and 4-Car Platoon), aerial drone control (Quadcopter), and dynamic balancing (CartPole and Pendulum). Table 1 summarizes each control system along with its associated safety properties. Specifically,
  • Quadcopter: This system stabilizes a quadcopter in flight by ensuring the pitch angle $\eta_{1}$ and yaw angle $\eta_{2}$ remain within $\pm\pi/2$ radians.
  • CartPole-v1 and CartPole-v2: Both systems involve a pole mounted on a cart that moves along a track. Safety is defined by the following constraints: the cart position $|\delta|<0.3$ m, cart velocity $|\delta^{\prime}|<0.5$ m/s, pole angle $|\eta|<30^{\circ}$, and pole angular velocity $|\eta^{\prime}|<0.5$ rad/s. The two variants differ in the complexity of neural policies: CartPole-v1 uses a neural policy with 0.7M neurons, while CartPole-v2 employs a larger policy with 7M neurons.
  • Pendulum-v1 and Pendulum-v2: These two systems aim to balance a pendulum in an upright position. Safety is maintained by restricting the angle $\eta$ to within $30^{\circ}$ and angular velocity $\omega$ to within $30^{\circ}$/s. The main difference lies in the pendulum length: 1.0 m for Pendulum-v1 and 1.5 m for Pendulum-v2.
  • Self-Driving-v1 and Self-Driving-v2: These systems simulate autonomous driving at constant speed, where safety requires the car’s heading angle $\eta$ to stay within $90^{\circ}$, and the lateral distance $d$ from the centerline to be less than 2.0 m for v1 and 1.5 m for v2, ensuring the vehicle does not veer into roadside canals.
  • 4-Car Platoon: This system models four vehicles maintaining safe distances. Each variable in the above specification is the distance between two consecutive vehicles in the platoon, and safety is assured if they are below certain thresholds.